OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. Answer y - 2 = 4(x + 6) y - 6 = 4(x + 2) y + 6 = 4(x - 2) y + 2 = 4(x - 6)

OpenStudy (anonymous):

Expand the brackets

OpenStudy (anonymous):

and simplify

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

NOOOPE

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y-2=4x-0.83 y-4x=-0.83+2 y-4x=1.17

OpenStudy (anonymous):

thats the first one

OpenStudy (anonymous):

Now equation of a line always carries the form y=mx+c where m is the gradient and c is the y intercept Now if we take the first simplified equation it is going to be y=4x+1.17

OpenStudy (anonymous):

it has the slope 4 or gradient 4

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

Thank You sooo much =)

OpenStudy (anonymous):

Sure for practice sake do the remaining ones and indicate the slope and y intercept

OpenStudy (anonymous):

ok